\begin{thebibliography}{99}                                           

\bibitem{aly} E. H. Aly, L. Elliott \& D. B. Ingham, Mixed convection boundary-layer flows over a vertical surface embedded in a porous medium, Eur. J. Mech. B Fluids 22 (2003), pp. 529-543.

\bibitem{banks1}W. H. H. Banks, Similarity solutions of the boundary layer equations
for a stretching wall, J. de M\'echan. Th\'eor. et  Appl. 2 (1983), pp. 375-392.

\bibitem{banks}W. H. H. Banks, M. B. Zaturska, Eigensolutions in boundary layer flow
adjacent to a stretching wall, IMA J. Appl. Math. 36 (1986), pp. 263-273.                                                     

\bibitem{brighi03}Z. Belhachmi, B. Brighi \& K. Taous, On the concave
solutions of the Blasius equation, Acta Math. Univ. Comenianae, Vol. LXIX, 2 (2000), 
pp. 199-214.

\bibitem{brighicr}Z. Belhachmi, B. Brighi \& K. Taous, Solutions similaires pour un 
probl\`eme de couche limite en milieux poreux, C. R. M\'ecanique 328 (2000), 
pp. 407-410.

\bibitem {brighi02}Z. Belhachmi, B. Brighi \& K. Taous, On a family of
differential equations for boundary layer approximations in porous media,
Euro. Jnl of Applied Mathematics, Vol. 12, 4, Cambridge University Press (2001),
pp. 513-528.

\bibitem {brighi04}Z. Belhachmi, B. Brighi, J. M. Sac-Epee \& K. Taous,
Numerical simulations of free convection about a vertical flat plate embedded
in a porous medium, Computational Geosciences, Vol. 7 (2003), pp. 137-166.

\bibitem{bla}H. Blasius, Grenzschichten in Fl\"ussigkeiten mit kleiner Reibung,
Z. Math. Phys. 56 (1908), pp. 1-37.

\bibitem {brighi01}B. Brighi, On a similarity boundary layer equation,
Zeitschrift f\"ur Analysis und ihre Anwendungen, Vol. 21, 4 (2002), pp. 931-948.

\bibitem {heat_flux}B. Brighi, J.-D. Hoernel, On similarity solutions for boundary layer flows with prescribed heat flux. Mathematical Methods in the Applied Sciences, Vol. 28, 4 (2005), pp. 479-503.

\bibitem{gaeta} B. Brighi, J.-D. Hoernel, Recent advances on similarity solutions arising during free convection, Progress in Nonlinear Differential Equations and Their Applications, Vol. 63, pp. 83-92, Birkh\"auser Verlag Basel/Switzerland, 2005.

\bibitem{equiv}B. Brighi, J.-D. Hoernel, Asymptotic behavior of the unbounded 
solutions of some boundary layer equation, Archiv der Mathematik, Vol. 85, 2 (2005), pp. 161-166.


\bibitem {aml}B. Brighi, J.-D. Hoernel, On the concave and convex solutions of mixed convection boundary layer approximation in a porous medium, Applied Mathematics Letters, Vol. 19, 1 (2006), pp. 69-74.

\bibitem {brighisari}B. Brighi, T. Sari, Blowing-up coordinates for a
similarity boundary layer equation. Discrete and Continuous Dynamical Systems 
(Serie A), Vol. 12, 5 (2005), pp. 929-948.

\bibitem {pop1}M. A. Chaudhary, J.H. Merkin \& I. Pop, Similarity solutions in
free convection boundary-layer flows adjacent to vertical permeable surfaces
in porous media: I prescribed surface temperature, Eur. J. Mech. B-Fluids, 14 (1995), pp. 217-237.

\bibitem {pop}M. A. Chaudhary, J.H. Merkin \& I. Pop, Similarity solutions in
free convection boundary-layer flows adjacent to vertical permeable surfaces
in porous media: II prescribed surface heat flux, Heat and Mass Transfer 30,
Springer-Verlag (1995), pp. 341-347.

\bibitem{cheng}P. Cheng, W. J. Minkowycz, Free-convection about a vertical flat
plate embedded in a porous medium with application to heat transfer from a dike,
J. Geophys. Res. 82 (14) (1977), pp. 2040-2044.

\bibitem{coppel} W. A. Coppel, On a differential equation of boundary layer theory, {\it Phil. Trans. Roy. Soc. London, Ser A} {\bf  253}, pp. 101-136 (1960).

\bibitem{crane}L. E. Crane, Flow past a stretching plane, Z. Angew. Math. Phys.
21 (1970), pp. 645-647.

\bibitem{ene} E. I. Ene, D. Poli\v{s}evski, Thermal flow in porous media, D. Reidel
Publishing Company, Dordrecht, 1987.

\bibitem{falk}V. M. Falkner, S. W. Skan, Solutions of the boundary layer equations, Phil.
Mag., 7/12 (1931), pp. 865-896.

%\bibitem{fed}M. V. Fedoryuk, Asymptotic Analysis, Srpinger-Verlag, 1993.

\bibitem{guedda} M. Guedda, Nonuniqueness of solutions to differential equations for boundary layer approximations in porous media, C. R. M\'ecanique, 330 (2002), pp. 279-283.

\bibitem{guedda1}M. Guedda, Similarity solutions of differential equations for boundary
layer approximations in porous media. To appear in ZAMP.

\bibitem{guedda2} M. Guedda, Multiple solutions of mixed convection boundary-layer approximations in a porous medium. To appear in Applied Mathematics Letters.

\bibitem{gup}P. S. Gupta, A. S. Gupta, Heat an mass transfer on a stretching sheet
with suction or blowing, Can. J. Chem. Eng. 55 (1977), pp. 744-746.

\bibitem {hart}P. Hartmann, Ordinary Differential Equations. Wiley, New-York (1964).

\bibitem{hast}S. P. Hastings, W. C. Troy, Oscillating solutions of the
Falkner-Skan equations for negative $\beta$, SIAM J. Math. Anal. 18(2) (1987), 
pp. 422-429.

\bibitem{hast1}S. P. Hastings, W. C. Troy, Oscillating solutions of the
Falkner-Skan equations for positive $\beta$, J. Diff. Equations 71(1) (1988), 
pp. 123-144.

\bibitem{ing}D. B. Ingham, S. N. Brown, Flow past a suddenly heated vertical plate
in a porous medium, J. Proc. R. Soc. Lond. A 403 (1986), pp. 51-80.

\bibitem{ish1}N. Ishimura, T. K. Ushijima, An elementary approach to the analysis of exact solutions for the Navier-Stokes stagnation flows with slips, Arch. Math. 82 (2004), pp. 432-441.

%\bibitem{Ish} N. Ishimura, S. Matsui, On blowing up solutions of the Blasius equation,
%Discrete and Continuous Dynamical Systems, Vol. 9 (2003), pp. 985-992.

%\bibitem{kam}E. Kamke, Zur theorie der Systeme gew\"ohnlicher Differentialgleichungen
%II, Acta Math. 58 (1932), pp. 57-85.

%\bibitem{kay}W. M. Kays, M. E. Crawford, Convective heat and mass transfer,
%Mc Graw Hill, Inc., 1993.

\bibitem{mag} E. Magyari, B. Keller, Exact solutions for self-similar boundary-layer flows induced by permeable stretching wall. Eur. J. Mech. B-Fluids 19 (2000), pp. 109-122.

\bibitem{mag1}E. Magyari, I. Pop, B. Keller, The 'missing' self-similar free convection
boundary-layer flow over a vertical permeable surface in a porous medium, Transport in
Porous Media 46 (2002), pp. 91-102.

\bibitem{merk} J. H. Merkin, G. Zhang, On the similarity solutions for free convection in
a saturated porous medium adjacent to impermeable horizontal surfaces, W\"arme und
Stoff\"ubertr., 25 (1990), pp.179-184.

\bibitem{mof}H. K. Moffatt, High-frequency excitation of liquid metal systems, IUTAM Symposium: Metallurgical Application of Magnetohydrodynamics, (1982) Cambridge.

\bibitem{nazar} R. Nazar, N. Amin \& I. Pop, Unsteady mixed convection boundary-layer flow near the stagnation point on a vertical surface in a porous medium, {\it Int. J. Heat Mass Transfer} {\bf 47}, pp. 2681-2688 (2004).

%\bibitem{ole}O. A. Oleinik, V. N. Samokhin, Mathematical models in boundary layer
%theory, Chapman and Hall / CRC, Boca Raton, 1999.

%\bibitem{pran}L. L. Prandtl, \"Uber Fl\"ussigkeitsbewegungen bei sehr kleiner
%Reihung, Verh. d. III, Int. Math. Kongr., pp. 484-494, Heidelberg 1904, Leipzig:
%Teubner 1905.

%\bibitem{pran1}L. L. Prandtl, Mechanics of viscous fluids, In: W. F. Daniel, editor,
%Aerodynamics theory (1935), pp. 34-208.

%\bibitem{rose}L. Rosenhead, Laminar boundary layers, Oxford University Press, 1963.

%\bibitem{sce}G. V. \v{S}\v{c}erbina, On a boundary problem for an equation of Blasius,
%Sov. Math., Dokl. 2 (1961), pp. 1219-1222.

%\bibitem{stu}?? J. T. Stuart, Double boundary layers in oscillatory viscous flow, J. Fluid. Mech. 24 (1966), pp. 673-687. 

%\bibitem{tam1}K. K. Tam, An elementary proof for the Blasius equation, Z. Angew.
%Math. Mech. 51 (1971), pp. 318-319.

%\bibitem{Tam}K. K. Tam, On the existence and parameter dependance of similarity
%solutions for free convection on a vertical plate, Studies in Applied Mathematics 88 (1993), 
%pp. 141-152.

\bibitem{utz}W. R. Utz, Existence of solutions of a generalized Blasius equation, J.
Math. Anal. Appl. 66 (1978), pp. 55-59.

%\bibitem{walt}W. Walter, Differential and Integral Inequalities, Springer-Verlag, Berlin
%Heidelberg New-York, 1970.

\bibitem{wang}J. Wang, W. Gao \& Z. Zhang, Singular nonlinear boundary value problems arising in boundary layer theory, J. Math. Anal. Appl. 233 (1999), pp. 246-256.

%\bibitem{wat}J. Watson, On the existence of solutions for a class of rotating-disc flows
%and the convergence of a successive approximation scheme, J. Inst. Math. Appl.,
%1 (1965), pp. 348-371.

%\bibitem{weyl}H. Weyl, On the differential equations of the simplest boundary-layer
%problem, Ann. Math. 43 (1942), pp. 381-407.

%\bibitem{wood}R. A. Wooding, Convection in a saturated porous medium at large Rayleigh number or Peclet number, J. Fluid. Mech., 15 (1963), pp. 527-544.

\bibitem{yang} G.C. Yang, Existence of solutions to the third-order nonlinear differential
equations arising in boundary layer theory, Appl. Math. Lett. 16 (2003) (6),
pp. 827-832. 

\bibitem{yang2} G.C. Yang, A note on $f'''+ff''+\l(1 - f'^2)=0$ with $\l\in (-1/2, 0)$
arising in boundary layer theory, Appl. Math. Lett. 17 (2004) (11), pp. 1261-1265. 

\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage{float}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{graphicx}%
\renewcommand{\baselinestretch}{1}
\newtheorem{theorem}{Theorem}

\newtheorem{lemma}{Lemma}

\newtheorem{proposition}{Proposition}
\newtheorem{remark}{Remark}

\newenvironment{proof}[1][Proof]{\noindent\textbf{#1.} }{\ \rule{0.5em}{0.5em}}

\let \a = \alpha
\let \b = \beta
\let \l = \lambda
\let \g = \gamma

\def \R {{\mathbb R}}

\begin{document}

\title{\bf On a general similarity boundary layer equation}
\author{BERNARD BRIGHI$^a$  and JEAN-DAVID HOERNEL$^{b,*}$ }
\date{}
\maketitle

\begin{center}
Universit\'e de Haute-Alsace 
\vskip 0,1cm
Laboratoire de Math\'ematiques, Informatique et Applications
\vskip 0,1cm
4 rue des fr\`eres Lumi\`ere, 68093 MULHOUSE (France)
\end{center}

\begin{abstract}
In this paper we are concerned with the solutions of the differential equation $f^{\prime\prime\prime}+ff^{\prime\prime}+g(f^{\prime})=0$ on $[0,\infty)$, satisfying the boundary conditions $f(0)=\alpha$, $f'(0)=\beta\geq 0$, $f'(\infty)=\l$, and where $g$ is some given continuous function.
This general boundary value problem includes the Falkner-Skan case, and can be applied, for example, to free or mixed convection in porous medium, or flow adjacent to stretching walls in the context of boundary layer approximation.
Under some assumptions on the function $g$, we prove existence and uniqueness of a concave or a convex solution. We also give some results about nonexistence and asymptotic behaviour of the solution.
\end{abstract}

\noindent AMS 2000 Subject Classification: 34B15; 34C1; 76D10.

\bigskip
\noindent Key words and phrases: Boundary layer, similarity solution, third order nonlinear differential equation,  boundary value problem, Falkner-Skan, free convection, mixed convection.

\renewcommand{\baselinestretch}{1}
\footnotetext{\hspace{-0.8cm}$^*$Corresponding author.}
\footnotetext{\hspace{-0.8cm}E-mail addresses: bernard.brighi@uha.fr, j-d.hoernel@wanadoo.fr}

\section{Introduction}
We consider the following third order non-linear autonomous differential
equation
\begin{equation}
f^{\prime\prime\prime}+ff^{\prime\prime}+g(f^{\prime})=0 \label{eq}
\end{equation}
with the boundary conditions
\begin{eqnarray}
f(0)=\alpha, \label{c1} \\
f'(0)=\beta, \label{c2} \\
f'(\infty)=\l \label{c3}
\end{eqnarray}
where $\a\in \R$, $\b\in \R_+$, $\l \in \R$ and $f'(\infty):=\underset{t\rightarrow\infty}{\lim} f'(t)$. We also assume that the given function $g$ is locally Lipschitz on some interval $J$ containing $\b$ and $\l$.

\newpage


In the litterature, problem (\ref{eq})-(\ref{c3}) with suitable $g$ and $\l$ arises in many fields of application such as free or mixed convection in a fluid saturated porous medium near a semi or double infinite wall in the framework of boundary layer approximation, high frequency excitation of liquid metal, stretching walls,...  

The problems of free convection, stretching walls and high frequency excitation of liquid metal corresponds to the function $g$ given by $g(x)=\frac{2m}{m+1}(-x)x$ and to $\l=0$. There are many papers in connection with those such as \cite{banks1}, \cite{banks},  \cite{brighi04},  \cite{pop1}, \cite{pop}, \cite{cheng}, \cite{crane}, \cite{ene}, \cite{gup}, \cite{ing}, \cite{mag}, \cite{mag1}, \cite{merk}, \cite{mof} for the physical point of view or numerical computations,  \cite{brighicr}, \cite{brighi02}, \cite{brighi01},  \cite{heat_flux}, \cite{equiv}, \cite{brighisari}, \cite{guedda}, \cite{guedda1} for the mathematical analysis and \cite{gaeta} for a survey.

The Falkner-Skan equation, arising in the study of two dimensional flow of a slightly viscous incompressible fluid past a wedge of angle $\pi m$ under the assumptions of boundary-layer theory, is obtained for $g(x)=m(1-x)(1+x)$ and $\l=1$. This famous equation has been widely studied, see for example \cite{coppel}, \cite{falk}, \cite{hart}, \cite{hast}, \cite{hast1} and the references therein for a survey, and \cite{ish1}, \cite{wang}, \cite{yang}, \cite{yang2} for more recent investigations.

The mixed convection case corresponds to $g(x)=\frac{2m}{m+1}(1-x)x$ and  $\l=1$. This problems appears recently in \cite{aly} and \cite{nazar}. Some first theoretical results about this equation can be found in \cite{aml} and \cite{guedda2}.

The Blasius problem, corresponding to $g=0$, is a particular case of all the previous situations and the first historic case in which an equation of the form (\ref{eq}) appears.
 This well known problem, that arises in \cite{bla} at the begining of the previous century, has been studied in a lot of papers. For more details, we refer to \cite{brighi03}, \cite{coppel} and \cite{hart} and the references therein.


Finally, let us notice that a first generalization of some of the previous equations can be found in \cite{utz}. The author considers problem (\ref{eq})-(\ref{c3}) with $\a\geq 0$, $\l\geq \b>0$ and functions $g$ such that $g(x)=\hat g(x^2)$ where $\hat g$ is assumed to be positive and monotone decreasing on $[\b,\l)$ and $\hat g(\l^2)=0$. Under these hypotheses, he proves that there exists one and only one convex solution of this problem.

\begin{remark}
Let $a>0$, and consider the differential equation $u'''+auu''+h(u')=0$ where $h$ is some given function $h$. By setting $u(t)=\frac{1}{\sqrt{a}}f(\sqrt{a}t)$ then the equation in $u$ reduces to $f'''+ff''+g(f')=0$ with $g(x)=\frac{1}{a}h(x)$. 


\end{remark}

\section{Preliminary results}

First of all, let us remark that if $f$ satisfies equation (\ref{eq}) on some interval $I$ and if we denote by $F$ any anti-derivative of $f$ on $I$, then we have
\begin{equation}
\left (f''e^F\right )'=-g(f')e^F.\label{exp}
\end{equation}
This, in particular implies that the concavity of $f$ is related to the changes of sign of $g$.

We will need some Lemmas concerning the solutions of (\ref{eq}).
\begin{lemma}\label{vanish}
If $g(\mu)=0$ and if $f$ is a solution of {\rm (\ref{eq})} on some interval $I$ such that there exists a point $t_0\in I$ verifying $f''(t_0)=0$ and $f'(t_0)=\mu$, then $f''(t)=0$ for every $t\in I$.
\end{lemma}
\begin{proof}
Let $f$ be a solution of (\ref{eq}) on $I$ such that $f''(t_0)=0$ and Ê$f'(t_0)=\mu$ for some $t_0\in I$. Since the function $r(t)=\mu(t-t_0)+f(t_0)$ is a solution of (\ref{eq}) such that $r(t_0)=f(t_0)$, $r'(t_0)=f'(t_0)$ and $r''(t_0)=f''(t_0)$, we get $r=f$ and $f''\equiv 0$.
\end{proof}

\begin{lemma}\label{f2} Let $f$ be a solution of $(\ref{eq})$ on some interval
$[t_0,\infty)$, such that $f'(t)\to l \in\R$ as $t\to\infty$. If moreover $f$ is of constant sign at infinity, then, we have
$$\lim_{t\to\infty}f''(t)=0.$$
\end{lemma}
\begin{proof} First of all, let us remark that since $f'(t)$ has a finite limit as
$t\to\infty$, then 
\begin{equation}
\liminf_{t\to\infty}f''(t)^2=0.\label{i}
\end{equation}
Multiplying $(\ref{eq})$ by $f''$ and integrating on $[t_0,t]$, we get 
\begin{equation}
\frac{1}{2}f''(t)^2-\frac{1}{2}f''(t_0)^2+\int_{t_0}^tf(s)f''(s)^2ds
+G(f'(t))-G(f'(t_0))=0,\label{ii}
\end{equation} 
where we denoted by $G$ any anti-derivative of $g$. 
As $f$ is of constant sign at infinity, it follows that the integral in $(\ref{ii})$, and
thus $f''(t)^2$ too, have limits as $t\to\infty$. From $(\ref{i})$ we get the result.
\end{proof}

\begin{remark} If $l \not=0$ we have $f(t)\sim l t$ as $t\to\infty$ and $f$ is of constant sign at infinity.
If $l=0$ and if $f$ is either concave or convex at infinity, then again $f$ is of constant sign
at infinity. From $(\ref{exp})$ it is the case, for example, if $g$ is of constant sign in a neighbourhood of $0$.
\end{remark}

The following  Lemma shows that to expect a solution of $(\ref{eq})$-$(\ref{c3})$, we must assume that the function $g$ vanishes at the point $\l$. 

\begin{lemma} \label{lim} Let $f$ be a solution of $(\ref{eq})$ on some interval $[t_0,\infty)$, such that $f'(t)\to l\in \R$ as $t\to\infty$. Then $g(l)=0$.
\end{lemma}
\begin{proof}
Let us suppose that $2c=-g(l)>0$. There exists $t_1>t_0$ such that $-g(f'(t))>c$ for $t>t_1$ and from (\ref{exp}) we have
$$\left ( f''e^{F}\right )'>ce^{F}$$
on $[t_1,\infty)$. This means that $f''$ cannot vanish more than once and thus $f$ is concave or convex at infinity.
\begin{itemize}
\item Assume now that $f$ is bounded. Then, it follows from $(\ref{eq})$ and Lemma \ref{f2} that $f'''(t)\to 2c$ as $t\to\infty$ and we have a contradiction.
\item Assume next that $f$ is unbounded. Then $\vert f(t)\vert\to\infty$ as $t\to\infty$ and thus there exists $t_2>t_1$ such that $f'/f>0$ on $[t_2,\infty)$. From $(\ref{eq})$ we get 
\begin{equation}
\frac{f'''f'}{f}+f''f'>c\frac{f'}{f} \quad \text{on} \quad [t_2,\infty).\label{iii}
\end{equation}
Integrating we obtain
$$\int_{t_2}^t\frac{f'''(s)f'(s)}{f(s)}ds+\frac{1}{2}f'(t)^2-\frac{1}{2}f'(t_2)^2\geq c\left(\ln\vert f(t)\vert-\ln\vert f(t_2)\vert\right).$$
It follows that 
\begin{equation}
\int_{t_2}^\infty\frac{f'''(s)f'(s)}{f(s)}ds=\infty.\label{iv}
\end{equation}
But, we have
\begin{align*}
\int_{t_2}^t\frac{f'''(s)f'(s)}{f(s)}ds= &\ \frac{f''(t)f'(t)}{f(t)}-\frac{f''(t_2)f'(t_2)}{f(t_2)} \cr
&\ -\int_{t_2}^t \frac{f''(s)^2}{f(s)}ds+\int_{t_2}^t \frac{f''(s)f'(s)^2}{f(s)^2}ds
\end{align*}
which leads to a contradiction with $(\ref{iv})$, since the integrals in the right hand side have
finite limits as $t\to\infty$. 

For $c<0$, same arguments give also a contradiction. The proof is now complete.
\end{itemize}
\vspace{-3mm}
\end{proof}

\begin{remark}
If $l \not = 0$ we can have a much simpler proof. Indeed, in this case we have $f'(t)\sim l$ and $f(t) \sim lt$ as $t \to \infty$. Integrating $(\ref{eq})$ on $[t_0,t]$ we get
\begin{align*}
f''(t)-f''(t_0)-f(t_0)f'(t_0)&=-f(t)f'(t)+\int_{t_0}^t f'(s)^2ds-\int_{t_0}^tg(f'(s))ds \cr
&=-l^2t(1+o(1))+l^2t(1+o(1))-g(l)t(1+o(1)) \cr
&=-g(l)t+o(t)
\end{align*}
which is a contradiction since $f''(t)\to 0$ as $t\to \infty$ by Lemma $\ref{f2}$.
\end{remark}

\begin{remark} Solution, for which first derivative does not have a finite limit, does exist. 
For example
\begin{itemize} \item For any $a\in\R^*$ and any $b\in\R$ the function $f$ defined by
$$f(t)=at^2+bt+\frac{b^2-1}{4a}$$
is a solution of $(\ref{eq})$ with $g(x)=\frac{1}{2}(1-x^2)$ and $f'(t)\to \pm\infty$ as $t\to \infty$ $($Falkner-Skan with $m=\frac{1}{2})$.
\item If $g(x)=-x^2+x+1$, then $f(t)=\sin t$ is a solution of $(\ref{eq})$ for which $f'$ does not have a limit at infinity.
\end{itemize}
\end{remark}


In order to get solutions of (\ref{eq})-(\ref{c3}) for given $\alpha\in\mathbb R$, $\beta \in\mathbb R_+$ and $\l\in\mathbb R$, we will consider the  initial value problem 
\begin{equation}
\left \{ \begin{array}
[c]{lll}
f'''+ff''+g(f')=0,\\ 
f(0) = \a,\\
f'(0) = \b,\\
f''(0) = \g
\end{array} \right . \label{initial}
\end{equation}
and use a shooting technique on the parameter $\gamma$. We will denote by $f_\gamma$ its solution and by $[0,T_\gamma)$ its right maximal interval of existence. 
Integrating (\ref{eq}) on $[0,t]$ for $0<t<T_\gamma$, we obtain the useful identity
\begin{equation}
f''_\gamma(t)=\g -f_\g(t)f'_\g(t)+\a\b+\int_0^t \left (f_\g'(s)^2-g(f_\g'(s))\right )ds. \label{i1}
\end{equation}

\begin{remark}
Looking at $(\ref{i1})$ we see that if we take $g(x)=x^2$ the integral in the right-hand side vanishes. Then, integrating on $[0,t]$ we obtain
$$f'(t)+\frac{1}{2}f(t)^2=(\g+\a\b)t+\frac{\a^2}{2}+\b$$
and choosing $\g=-\a\b$ we have that for $\b>-\a^2/2$ the function
$$f(t)=\frac{1}{-\frac{1}{2d}+\frac{\a+d}{2\a d-2d^2}e^{dt}}+d$$
with $d=\sqrt{\a^2+2\b}$ is a bounded solution of the problem {\rm (\ref{eq})-(\ref{c3})} for $\l=0$.
\end{remark}

\begin{remark}
Let us take a look at the case $\beta=\l>0$. If $g(\l)=0$, then the function $f_0(t)=\l t+\alpha$ is  a solution of $(\ref{eq})$-$(\ref{c3})$. Without additional hypotheses on $g$, we cannot say anythings about uniqueness. However, if we assume, for example, that $g(x)<0$ for $x>\l$ and $g(x)>0$ for $x<\l$, then $f_0$ is the unique solution of $(\ref{eq})$-$(\ref{c3})$. Indeed, let $f_\g$ be another solution of $(\ref{eq})$-$(\ref{c3})$ such that $\gamma>0$. Since $f_\g'(0)=f_\g'(\infty)=\l$, there exists $t_0>0$ such that $f_\g'(t_0)>\l$, $f_\g''(t_0)=0$ and $f_\g'''(t_0)\leq 0$. From $(\ref{eq})$ we obtain $f_\g'''(t_0)=-g(f_\g'(t_0))>0$ and thus a contradiction. 
If $\gamma<0$, the same approach leads again to a contradiction.
\end{remark}

In the following, we will focus first on the concave solutions and next on the convex solutions of (\ref{eq})-(\ref{c3}) with functions $g$ such that $g(\l)=0$. As seen in Lemma \ref{lim}, this hypothesis  is necessary to realize the condition $f'(t)\to \lambda$ as $t\to \infty$. In addition, we will assume that some condition on the sign of $g$ is satisfied between $\b$ and $\l$ in order to get existence and uniqueness of a concave solution (when $\l<\b$) or a convex solution (when $\l>\b$) of the problem (\ref{eq})-(\ref{c3}).  Such an assumption holds in the physical cases evoked in the introduction for the positive values of the parameter $m$.

When the proofs in the convex case are close to the ones of the concave case we will remove some details in order to shorten them.

\section{Concave solutions}
\begin{theorem}\label{concave}
Let $\a \in \mathbb R$ and $0\leq \l<\b$. If $g(x)<0$ for $x\in(\l,\b]$ and $g(\l)=0$, then the problem {\rm (\ref{eq})-(\ref{c3})} admits a unique concave solution.
\end{theorem}
\begin{proof}[Proof of existence] 
Let $f_\gamma$ be a solution of the initial value problem (\ref{initial}) with $\l<\b$ and $\gamma\leq 0$. As long as we have $f'_\gamma>\l$ and $f''_\gamma<0$, $f_\gamma$ exists. Because of Lemma \ref{vanish}, there are only three possibilities

\vspace{3mm}
(a) $f''_\gamma$ becomes positive from a point such that $f'_\gamma>\l,$

(b) $f'_\gamma$ takes the value $\l$ at some point for which $f''_\gamma<0,$

(c) we always have $f'_\gamma>\l$ and $f''_\gamma<0.$

\vspace{3mm}
\noindent As $f'_0(0)=\b>\l$, $f''_0(0)=0$ and $f'''_0(0)=-g(f'_0(0))=-g(\b)>0$ we have that $f'_0(t)>\l$ and $f''_0(t)>0$ on some interval $[0,t_0)$. By continuity it follows that $f''_\gamma$ becomes positive at some point for which $f'_\gamma>\l$ for small values of $-\gamma$ and thus $f_\gamma$ is of type (a).

On the other hand, as long as $f''_\gamma(t)<0$ and $f'_\gamma(t)\geq \l$, we have $f_\gamma(t)\geq \a$, $f'_\gamma(t)\leq \b$ and, using (\ref{i1}) we obtain
\begin{align*}
f''_\gamma(t) &\leq \g +|\a|\b+\a\b+\b^2t-\int_0^t g(f_\g'(\xi))d\xi \cr
&\leq \g +|\a|\b+\a\b+(\b^2+C)t
\end{align*}
where $C=\max\{-g(x)~;~x\in[\l,\b]\}>0$. Integrating once again we have
$$\l \leq f'_\gamma(t) \leq \frac{\b^2+C}{2}t^2+(\g+\a\b+|\a|\b)t+\b:=P_\g(t).$$
Hence, for $-\gamma$ large enough, the equation $P_\g(t)=\l$ has two positive roots $t_0<t_1$, and therefore we have $f'_\g(t_0)=\l$ and $f''_\gamma(t)<0$ for $t\leq t_0$, and $f_\gamma$ is of type (b).

Defining $A=\left \{ \gamma<0\, ; \, \text{$f_\gamma$ is of type (a)}\right \}$ and
$B=\left \{ \gamma<0\, ; \, \text{$f_\gamma$ is of type (b)}\right \}$ we have that 
$A \neq \emptyset$, $B \neq \emptyset$ and $A \cap B = \emptyset$. Both $A$ and $B$ are open sets, so there exists a $\gamma_*<0$ such that the solution $f_{\gamma_*}$ of (\ref{initial}) is of type (c) and is defined on the whole interval $[0,\infty)$. For this solution we have that $f'_{\gamma_*}>\l$ and $f''_{\gamma_*}<0$ which implies that $f'_{\gamma_*} \rightarrow l \in[\l,\b)$ as $t\rightarrow \infty$. From Lemma \ref{lim} and the fact that $g<0$ on $(\l,\b]$ we get $l=\l$.
\end{proof}

\vspace{3mm}
\begin{proof}[Proof of uniqueness]
Let $f$ be a concave solution of (\ref{eq})-(\ref{c3}). As $f'$ is positive and strictly decreasing, we can define a function $v:(\l^2,\b^2]\rightarrow[\a,\infty)$ such that
$$\forall t\geq 0,\quad v(f'(t)^2)=f(t).$$
Setting $y=f'(t)^2$ leads to
\begin{equation}
f(t)=v(y),\quad f''(t)=\frac{1}{2v'(y)}\quad \text{and} \quad f'''(t)=-\frac{v''(y)\sqrt{y}}{2v'(y)^3}.\label{y1}
\end{equation}
Then, using (\ref{eq}) we obtain
\begin{equation}
\forall y\in (\l^2,\b^2],\quad v''(y)=\frac{v(y)v'(y)^2}{\sqrt{y}}+\frac{2v'(y)^3g(\sqrt{y})}{\sqrt{y}} \label{eq-v1}
\end{equation}
with
$$v(\b^2)=v(f'(0)^2)=\a \quad  \text{and}\quad v'(\b^2)=\frac{1}{2\g}<0.$$
Suppose now that there are two concave solutions $f_{1}$ and $f_{2}$ of 
(\ref{eq})-(\ref{c3}) with $f''_i(0)=\gamma_i<0$ $i\in\{1,2\}$ and $\gamma_1>\gamma_2$. They gives $v_{1},\, v_{2}$ solutions of equation (\ref{eq-v1}) defined on
$(\l^2,\b^2]$ such that, if $w=v_1-v_2$, we have
$$w(\b^2)=0 \quad \text{and} \quad w'(\b^2)=\frac{1}{2\gamma_1}-\frac{1}{2\gamma_2}<0.$$
If $w'$ vanishes, there exists an $x$ in $(\l^2,\b^2]$ such that $w'(x)=0$, $w''(x)\leq0$ and $w(x)>0$. But from (\ref{eq-v1}) we then obtain
$$w''(x)=\frac{v'_1(x)^2}{\sqrt{x}}w(x)>0$$
and this is a contradiction. Therefore, $w'<0$ and $w>0$ on $(\l^2,\b^2]$. Set now $V_i=\frac{1}{v'_i}$ for $i\in\{1,2\}$ and $W=V_1-V_2$ to obtain
$$W'(y)=-\frac{w(y)}{\sqrt{y}}-\frac{2w'(y)g(\sqrt{y})}{\sqrt{y}}<0.$$
But, using (\ref{y1}) we have $V_i(f'_i(t)^2)=2f_i''(t)$ and thanks to Lemma \ref{f2} we get $W(y) \rightarrow 0$ as $y\rightarrow \l^2$. Since $W$ is decreasing and $W(\b^2)=2(\g_1-\g_2)>0$ this is a contradiction.
\end{proof}

\begin{remark} If, in addition to the hypotheses of Theorem $\ref{concave}$, the function $g$ is assumed to be non increasing, then we can write a much simpler proof for the uniqueness result.
For that, let $f_1$ and $f_2$ be two concave solutions of {\rm (\ref{eq})-(\ref{c3})} and let $\gamma_i=f''_i(0)<0$, $i\in \left\{ 1,2 \right \}$ with $\gamma_1>\gamma_2$. Writing $u=f_1-f_2$, we have $u'(0)=0$, $u'(\infty)=0$ and $u''(0)>0.$
Hence $u'$ admits a positive local maximum at some $t_0>0$ such that $u'(t)>0$ for $t\in(0,t_0]$. As $u$ is increasing on $[0,t_0]$ and $u(0)=0$ we have $u(t_0)>0$. Then, from $(\ref{eq})$ and since $f'_i>\l$, $f''_i<0$ and $f''_1(t_0)=f''_2(t_0)$ we get
\begin{equation*}
u'''(t_0)=-f''_1(t_0)u(t_0)+g(f_2'(t_0))-g(f_1'(t_0))>0
\end{equation*}
because $f_1'(t_0)>f_2'(t_0)>\l$ and a contradiction with the fact that $u'''(t_0)\leq 0$.
\end{remark}

The following Proposition gives some informations about the behaviour at infinity of the concave solution of the problem (\ref{eq})-(\ref{c3}) obtained in Theorem \ref{concave}.

\begin{proposition} Let $\a\in \R$ and $0\leq \l<\b$. Let us assume that $g<0$ on $(\l,\b]$ and $g(\l)=0$, and let $f$ be the concave solution of $(\ref{eq})$-$(\ref{c3})$. Then, there exists a constant $\mu$ such that $\a<\mu<\sqrt{\a^2+2(\b-\l)}$ and
\begin{equation}
\lim_{t\to\infty}\{f(t)-(\lambda t+\mu)\}=0. \label{p1}
\end{equation}
Moreover, for all $t\geq 0$, one has $\l t +\a \leq f(t) \leq \l t +\mu$.
\end{proposition}
\begin{proof} 
Since $f$ is concave, then for all $t\geq 0$ we
have
$f'(t)\in(\lambda,\beta]$ and the function $t\mapsto f(t)-\lambda t$ is
increasing. Hence $f(t)-\lambda t\to\mu\in(\alpha,\infty]$ as $t\to\infty$. In addition, we have
$$\forall t\geq 0,\quad f'''(t)=-f(t)f''(t)-g(f'(t))\geq-f(t)f''(t)$$
and thus 
$$\forall t\geq 0,\quad {f'''(t)\over -f''(t)}\geq f(t)\geq f(t)-\l t.$$
If we assume that $\mu=\infty$, it follows that
$$\lim_{t\to\infty}{f'''(t)\over -f''(t)}=\infty.$$
Therefore, there exists $t_0\geq 0$ such that $f'''(t)\geq -f''(t)$ for $t\geq t_0$.
Then integrating twice and using Lemma \ref{f2} we get 
$$\forall t\geq t_0,\quad -f''(t)\geq-\lambda+f'(t)$$
and 
$$\forall t\geq t_0,\quad -f'(t)+f'(t_0)\geq-\lambda t+\lambda t_0+f(t)-f(t_0).$$
Since the left hand side is bounded, we get a contradiction. Therefore, $\mu<\infty$ and we have
$$\forall t> 0,~~~~~~\lambda t+\a<f(t)<\lambda t+\mu.$$
Finally, let us introduce the auxiliary nonnegative function
$$u(t)=f'(t)+{1\over 2}(f(t)-\lambda t)^2.$$
From (\ref{p1}), we see that $u$ is bounded. Moreover, we have
$$u''(t)=-g(f'(t))-\lambda tf''(t)+(f'(t)-\lambda)^2>0$$
and $u$ is convex. Therefore $u$ is decreasing and thus
$$\b+{\a^2\over 2}=u(0)>u(\infty)=\l+{\mu^2 \over 2}.$$
This completes the proof.
\end{proof}

\begin{remark}
If $\l=0$ the previous result means that the concave solution of {\rm (\ref{eq})-(\ref{c3})} is bounded.
\end{remark}

We have the following nonexistence result
\begin{theorem} \label{nexist} 
Let $\alpha\leq 0$ and $0\leq\lambda<\beta$. If $g$ is differentiable and if
\begin{equation}
\forall x\in[\lambda,\beta],\quad g(x)\geq x^2-\lambda x\quad \text{and} \quad
-\alpha+\max_{x\in[\lambda,\beta]}\{g(x)-x^2+\lambda x\}>0, \label{g1}
\end{equation}
then the problem $(\ref{eq})$-$(\ref{c3})$ does not admit concave solutions.
\end{theorem}
\begin{proof}
We follow an idea of \cite{wang}. Let $0\leq \l<\b$ and suppose that $f$ is a concave solution of (\ref{eq})-(\ref{c3}). As $f'$ is positive and strictly decreasing, we can define a negative function $v:(\l,\b]\rightarrow {\mathbb R}$ such that
$$\forall t\geq 0,\quad v(f'(t))=f''(t).$$
Setting $y=f'(t)$ we obtain
\begin{equation}
f''(t)=v(y), \quad f'''(t)=v(y)v'(y) \quad \text{and} \quad f^{(4)}(t)=v(y)^2v''(y)+v(y)v'(y)^2.
\label{vy}
\end{equation}
Derivating equation (\ref{eq}) leads to
$$f^{(4)}+f'f''+ff'''+g'(f')f''=0$$
and as
$$f=\frac{-f'''-g(f')}{f''}$$
we get
\begin{equation}
v''(y)=-\frac{y}{v(y)}-\left ( \frac{g(y)}{v(y)} \right )'. \label{eqv}
\end{equation}
Integrating (\ref{eqv}) on $[z,z_1]$ with $\l\leq z \leq z_1$ leads to
\begin{equation}
v'(z_1)-v'(z)=-\int_z^{z_1} \frac{y}{v(y)}dy-\frac{g(z_1)}{v(z_1)}+\frac{g(z)}{v(z)}. \label{iv1}
\end{equation}
Using equation (\ref{eq}) and (\ref{vy}), equality (\ref{iv1}) becomes 
$$-f(s_1)-v'(z)=-\int_z^{z_1} \frac{y}{v(y)}dy+\frac{g(z)}{v(z)}$$
with $s_1$ such that $f'(s_1)=z_1$.

Integrating on $[\l,x]$ with $\l\leq x \leq z_1$, and since $v(\l)=f''(\infty)=0$ by Lemma \ref{f2}, we get 
\begin{align*}
-f(s_1)(x-\l)-v(x)&=-\int_\l^x\left ( \int_z^{z_1} \frac{y}{v(y)}dy \right )dz+\int_\l^x\frac{g(z)}{v(z)}dz \cr
&=-\left [ z \int_z^{z_1} \frac{y}{v(y)}dy \right ]_\l^x+ \int_\l^x \frac{-z^2+g(z)}{v(z)}dz \cr
&=-x \int_x^{z_1} \frac{y}{v(y)}dy+\l \int_\l^{z_1} \frac{z}{v(z)}dz + \int_\l^x \frac{-z^2+g(z)}{v(z)}dz
\end{align*}
and taking $x={z_1}$ and $z_1\to \b$ we derive
\begin{equation}
-v(\b)= \int_\l^\b \frac{\l z-z^2+g(z)}{v(z)}dz+\a(\b-\l). \label{int1}
\end{equation}
Since $v(\beta)\leq 0$, the right hand side of $(\ref{int1})$ must be nonnegative,
but this cannot be the case if $(\ref{g1})$ holds.
\end{proof}


\begin{remark}Using the previous Theorem we can recover the following nonexistence result
\begin{itemize}
\item for free convection $($i.e. $g(x)=-\frac{2m}{m+1}x^2$ and $\l=0$$)$ there is no concave solutions for $-1<m\leq -\frac{1}{3}$ when $\a<0$, and $-1<m< -\frac{1}{3}$ when $\a=0$ $($see {\rm \cite{brighicr}}, {\rm \cite{brighi02}}, {\rm \cite{brighi01}} and {\rm \cite{brighisari}}$)$,
\end{itemize}
and obtain the new results
\begin{itemize}
\item for the Falkner-Skan case $($i.e. $g(x)=m(1-x)(1+x)$ and $\l=1$$)$ there is no concave solutions with $\a\leq 0$ and $\b>1$ for $m\leq -\frac{\b}{1+\b}$,
\item for mixed convection $($i.e. $g(x)=\frac{2m}{m+1}x(1-x)$ and $\l=1$$)$ there is no concave solutions with $\a\leq 0$ and $\b>1$ for $-1<m< -\frac{1}{3}$.
\end{itemize}
\end{remark}

\section{Convex solutions}
\begin{theorem} \label{convex}
Let $\a\in\mathbb R$ and $0\leq \b<\l$. If $g(x)>0$ for $x\in[\b,\l)$ and $g(\l)=0$, then the problem {\rm (\ref{eq})-(\ref{c3})} admits a unique convex solution. 
\end{theorem}
\begin{proof}[Proof of existence] Let $f_\gamma$ be a solution of the initial value problem (\ref{initial}) with $0\leq \b<\l$ and $\gamma\geq 0$. We notice that $f_\gamma$ exists as long as we have $f''_\g>0$ and $f'_\g<\l$. From Lemma \ref{vanish}, $f''_\gamma$ cannot vanish at a point where $f'_\gamma=\l$ and it follows that there are only three possibilities

\vspace{3mm}
(a) $f''_\gamma$ becomes negative from a point such that $f'_\gamma<\l,$

(b) $f'_\gamma$ takes the value $\l$ at some point for which $f''_\gamma>0,$

(c) we always have $\b\leq f'_\gamma<\l$ and $f''_\gamma>0.$
\vspace{3mm}

\noindent As $f'_0(0)=\b<\l$, $f''_0(0)=0$ and $f'''_0(0)=-g(\b)<0$, we have that $f_0$ is of type (a), and by continuity it must be so for $f_\gamma$ with $\gamma>0$ small enough.

On the other hand, as long as $f''_\gamma(t)>0$ and $f'_\gamma(t)\leq \l$, we have  $f_\gamma(t)\leq \l t+\a$, and (\ref{i1}) leads to
\begin{align*}
f''_\gamma(t) &\geq \g-f_\gamma(t)f'_\gamma(t)+\a\b-\int_0^tg(f'_\gamma(\xi))d\xi \cr
&\geq \g-(\l t+|\a|)\l+\a\b-\int_0^tg(f'_\gamma(\xi))d\xi \cr
&\geq \g-|\a|\l+\a\b-(\l^2+C)t 
\end{align*}
where $C=\max\{g(x)~;~x\in[\b,\l]\}>0$ and integrating once again we have
\begin{equation*}
\l\geq f'_\gamma(t) \geq -\frac{\l^2+C}{2}t^2+(\g-|\a|\l+\a\b)t+\b:=P_\g(t).\label{p}
\end{equation*}
Hence, for $\gamma$ large enough, the equation $P_\g(t)=\l$ has two positive roots $t_0<t_1$, and therefore, for such a $\g$, we have $f'_\gamma(t_0)=\l$ and $f''_\gamma(t)>0$ for $t\leq t_0$, and $f_\gamma$ is of type (b).

Defining $A=\left \{ \gamma>0\, ; \, \text{$f_\gamma$ is of type (a)}\right \}$ and
$B=\left \{ \gamma>0\, ; \, \text{$f_\gamma$ is of type (b)}\right \}$ we have that 
$A \neq \emptyset$, $B \neq \emptyset$ and $A \cap B = \emptyset$. Both $A$ and $B$ are open sets, so there exists a $\gamma_*>0$ such that the solution $f_{\gamma_*}$ of 
(\ref{initial}) is of type (c) and is defined on the whole interval $[0,\infty)$. For this solution we have that $0<f'_{\gamma_*}<\l$ and $f''_{\gamma_*}>0$ which implies that $f'_{\gamma_*} \rightarrow l \in (\b,\l]$ as $t\rightarrow \infty$. From Lemma \ref{lim} and the fact that $g>0$ on $[\b,\l)$ we get $l=\l$.
\end{proof}

\vspace{3mm}
\begin{proof}[Proof of uniqueness]
Let $f$ be a convex solution of (\ref{eq})-(\ref{c3}). As $f'$ and $f''$ are positive, we can define a function $v:[\b^2,\l^2)\rightarrow[\a,\infty)$ such that
$$\forall t\geq 0,\quad v(f'(t)^2)=f(t).$$
We have
\begin{equation}
\forall y\in [\b^2,\l^2),\quad v''(y)=\frac{v(y)v'(y)^2}{\sqrt{y}}+\frac{2v'(y)^3g(\sqrt{y})}{\sqrt{y}} \label{eq-v}
\end{equation}
and
$$v(\b^2)=v(f'(0)^2)=\a \quad \text{and}\quad v'(\b^2)=\frac{1}{2\g}>0.$$
Suppose now that there are two convex solutions $f_{1}$ and $f_{2}$ of 
(\ref{eq})-(\ref{c3}) with $f''_i(0)=\gamma_i>0$, $i\in\{1,2\}$ and $\gamma_1>\gamma_2$. They gives $v_{1},\, v_{2}$ solutions of equation (\ref{eq-v}) defined on
$[\b^2,\l^2)$ such that
for $w=v_{1}-v_{2}$, we have $w(\b^2)=0$ and $w'(\b^2)< 0$. 
If $w'$ vanishes, there exists an $x$ in $[\b^2,\l^2)$ such that $w'(x)=0$, $w''(x)\geq0$ and $w(x)<0$. But from (\ref{eq-v}) we then obtain $w''(x)<0$ and a contradiction. Therefore $w'<0$ and $w<0$ on $[\b^2,\l^2)$. Set now $V_i=\frac{1}{v'_i}$ for $i\in\{1,2\}$ and $W=V_1-V_2$. We have $W>0$ and using (\ref{eq-v}) we obtain $W'(y)>0.$
But, $W(\b^2)=2(\g_1-\g_2)>0$ and $W(y) \rightarrow 0$ as $y\rightarrow \l^2$, this contradict the fact that $W$ is increasing.
\end{proof}

\begin{proposition} Let $\a\in \R$ and $0\leq \b<\l$. Assume that $g>0$ on $[\b,\l)$ and $g(\l)=0$, and let $f$ be the convex solution of $(\ref{eq})$-$(\ref{c3})$. Then, there exists a constant $\mu>\a$ such that 
$$\lim_{t\to\infty}\{f(t)-(\lambda t+\mu)\}=0.$$
Moreover, for all $t\geq 0$, one has $\l t +\a \leq f(t) \leq \l t +\mu$.
\end{proposition}
\begin{proof} 
Since $f$ is convex, for all $t\geq 0$ we have
$f'(t)\in[\beta,\lambda)$. Then the function $t\mapsto f(t)-\lambda t$ is decreasing
and thus $f(t)-\lambda t\to\mu\in[-\infty,\alpha)$ as $t\to\infty$. On the other
hand, we have
$$\forall t\geq 0, \quad f'''(t)=-f(t)f''(t)-g(f'(t))\leq-f(t)f''(t)$$
and since $f(t)\to\infty$ as $t\to\infty$,
there exists $t_0\geq 0$ such that $f'''(t)\leq -f''(t)$ for $t\geq t_0$.
Then integrating twice and using Lemma \ref{f2} we get 
$$\forall t\geq t_0,\quad -f''(t)\leq-\lambda+f'(t)$$
and 
$$\forall t\geq t_0,\quad -f'(t)+f'(t_0)\leq-\lambda t+\lambda t_0+f(t)-f(t_0).$$
Since the left hand side is bounded, we necessarily get $\mu>-\infty$, and then we have $\l t +\mu \leq f(t) \leq \l t +\a$ for all $t\geq 0$.
\end{proof}

\bigskip
Let us finish this section with the following nonexistence result.
\begin{theorem}
Let $\alpha\leq 0$ and $0\leq\b<\l$. If $g$ is differentiable and if
\begin{equation}
\forall x\in[\b,\l],\quad g(x)\leq x^2-\lambda x\quad \text{and} \quad
-\alpha+\max_{x\in[\b,\l]}\{x^2-\lambda x-g(x)\}>0, \label{g2}
\end{equation}
then the problem $(\ref{eq})$-$(\ref{c3})$ does not admit convex solutions
\end{theorem}
\begin{proof}
Let $0\leq \b<\l$ and suppose that $f$ is a convex solution of (\ref{eq})-(\ref{c3}). As $f'$ and $f''$ are positive, we can define a positive function $v:[\b,\l)\rightarrow {\mathbb R}$ such that
$$\forall t\geq 0,\quad v(f'(t))=f''(t).$$
Following the method used in the proof of Theorem \ref{nexist}, we see that this function $v$ satisfies (\ref{eqv}) and 
\begin{equation}
-v(\b)= \int_\b^\l \frac{y^2-\l y-g(y)}{v(y)}dy-\a(\l-\b). \label{eqa}
\end{equation}
Since $v(\beta)\geq 0$, the right hand side of $(\ref{eqa})$ must be nonpositive,
but this cannot be the case if $(\ref{g2})$ holds.
\end{proof}
\bigskip

\begin{remark}Using the previous Theorem we can recover the following nonexistence result
\begin{itemize}
\item for the Falkner-Skan case $($i.e. $g(x)=m(1-x)(1+x)$ and $\l=1$$)$ there is no convex solutions for $m\leq -\frac{1}{2}$ $($see {\rm \cite{wang}}$)$,
\end{itemize}
and obtain the new result
\begin{itemize}
\item for mixed convection $($i.e. $g(x)=\frac{2m}{m+1}x(1-x)$ and $\l=1$$)$ there is no convex solutions for $-1<m\leq -\frac{1}{3}$.
\end{itemize}
\end{remark}


\section{Conclusion}
In this paper we have obtained existence, uniqueness and nonexistence results for the concave or convex solutions of a general boundary value problem arising in many fields of application under some reasonable hypotheses. 
All these hypotheses are verified in important physical cases in the framework of boundary layer approximations such as free or mixed convection
, flow adjacent to stretching walls, high frequency excitation of liquid metal and two dimensional flow of a slightly viscous incompressible fluid past a wedge. 


All our results hold for $\b$, $\l$ and $g$ such that the function $g$ vanishes at $\l$ but does not vanish between $\b$ and $\l$. 
Of course, under the same hypotheses, 
solutions whose concavity changes may exist as it can be seen in \cite{brighisari} for the case of free convection and in \cite{hast1} for the Falkner-Skan problem when  the parameter $m$ is positive.

In addition, if we assume that the sign of $g$ is the opposite of the one that we have in the Theorems \ref{concave} and \ref{convex}, then we can have multiple concave or convex solutions, as it is the case for free or mixed convection and for the Falkner-Skan problem when  the parameter $m$ is negative. See for example \cite{brighi01}, \cite{gaeta}, \cite{guedda}, \cite{guedda2} and \cite{hart}.



\end{document}